OpenStudy (anonymous):

Use the limit process to find the area of the region between the graph of y=x^2+3 and the x-axis over the interval [0,2]? thanks for any brave respones

OpenStudy (anonymous):

they give this triangle symbol which = 2/n

OpenStudy (zarkon):

\[\Delta x=\frac{2}{n}\]

OpenStudy (anonymous):

ok yeah thats the triangle but theres no x by it, thats why its confusing. Cant be delta x

OpenStudy (phi):

I would assume they mean delta x (width of each rectangle)

OpenStudy (zarkon):

\[\lim_{n\to\infty}\sum_{i=1}^{n}f(x_{i}^{*})\Delta x\]

OpenStudy (anonymous):

ok that makes sense with the delta x, thanks phi.... Zarkon ok so im assuming thats the formula. How would I take my next step? and thanks guys i really preciate this

OpenStudy (zarkon):

are you using left or right hand endpoints...or something else. (I'm assuming you are not using arbitrary partitions.)

OpenStudy (zarkon):

I would use \[x_{i}^*=a+i\Delta x\]

OpenStudy (anonymous):

The problem only gave me what i said above, I nerver even heard of using arbitrary partitions

OpenStudy (anonymous):

ok ill give it a try

OpenStudy (zarkon):

also I assume you meant y=x^2+3

OpenStudy (anonymous):

wow, yeah i did. My bad